OpenStudy (artfis):

How do you solve for x? square root of 3x + 11 = square root of 6x + 3

OpenStudy (artfis):

\[\sqrt{3x+11} = \sqrt{6x+3}\]

OpenStudy (mathstudent55):

Square both sides to get rid of square roots.

OpenStudy (mathstudent55):

Then solve like you'd solve a regular equation. Finally, you must check all answers because squaring equations may introduce extraneous solutions.

OpenStudy (mathstudent55):